Supose the.market for grass seed can be expressed as,follow as demand 100-2p,supply 3p at the equilibriium ,calculate the price elasticities of supply and demand ?what would happen with quantity demand if the price reach 30 and price reach to 50?
1
Expert's answer
2020-04-17T09:43:26-0400

With given supply and demand curve the equilibrium price and quantity are P=20. Q=60.

Point elasticity formula: 

"E = \frac{dQ}{dP}\times\frac{P}{Q}."

So, "Ed = -2\times\frac{20}{60}."

"Es = 3\times\frac{20}{60}=1."

The price elasticity of demand is -2/3, the price elasticity of supply is 1.

Demand curve "Q=100-2P" , so if the price is P=20, than the demand is "Q(30) = 100-2\times30=40."

If the price is 50, that the demand is "Q(50) = 100-2\times50=0."